A house wife find that 5vcans of condensed milk and 3 jars of instant coffee cost $27 while 12 cans of condensed milk and jars of instant coffee cost $49.40.find the total cost for 7 cans of condensed milk and 2 jars of instant coffee?
The bill for 6 cups of coffee and 7 cups of tea is $5.90 while the bill for 18 cups of coffee and 5 cup of tea is $9.70.find the total bill for 7 cups of coffee and 6 cups of tea?

A man bought 8 kiwi fruits and 7 pears for $4.10 while another man bought 4 kiwi fruits and 9 pears for $3.70.What is the cost of each kiwi fruit and each pear?

Answer by addingup(3677)

12 cans of condensed milk and how many jars of instant coffee?
====================================================================
6 cups of coffee and 7 cups of tea $5.90
18 cups of coffee and 5 cup of tea $9.70
find 7 cups of coffee and 6 cups of tea
:
6x + 7y = 5.90 multiply all times 3:
18x + 21y = 17.70 (1)
18x + 5y = 9.70 subtract this equation from (1):
-----------------------------------------------
18x + 21y = 17.70
-
18x + 5y = 9.70
----------------
0x + 16y = 8
16y = 8
y = 0.50 tea is 0.50 per cup
6x + 7y = 5.90
6x + 7(0.50) = 5.90
6x + 3.50 = 5.90
6x = 2.40
x = 0.40 coffee is 0.40 per cup
Now you find the cost of 7 cups of coffee and 6 cups of tea. Use your calculator.
The process is the same for all 3 problems, follow my work to solve the other two.
Happy learning,
John